# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 6.2 Randomized Algorithms for Tucker Decompositions with Orthogonal Factors
In particular, a CountSketch operator $\bm{\mathsf{S}}$ can be rapidly applied to a matrix $\bm{\mathsf{A}}$ in $O(\textup{nnz}(\bm{\mathsf{A}}))$ operations, without explicitly forming $\bm{\mathsf{S}}$ or $\bm{\mathsf{A}}$. More specifically, if $\bm{\mathsf{S}}\in\mathbb{R}^{J\times I}$ and $\bm{\mathsf{A}}\in\mathbb{F}^{I\times L}$, typically with $J\ll I$, then we can form $\bm{\mathsf{Y}}=\bm{\mathsf{S}}\bm{\mathsf{A}}$ by hashing each row $i$ of $\bm{\mathsf{A}}$ with an integer $h$ sampled uniformly from $[J]$, assigning it a value $s$ of $\pm 1$ with equal probability, then summing the rows with the same hash value (i.e. adding $s\bm{\mathsf{A}}_{i,:}$ to the $h$th row of $\bm{\mathsf{Y}}$ if $\bm{\mathsf{A}}_{i,:}$ has hash value $h$).
@@
A TensorSketch operator is a specific type of CountSketch operator, which can be rapidly applied to Kronecker products, e.g., $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)}\in\mathbb{F}^{n^{d}\times k^{d}}$. (We again assume here that $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$.) If $\bm{\mathsf{T}}:\mathbb{R}^{n^{d}}\rightarrow\mathbb{R}^{\ell}$ is a TensorSketch operator with embedding dimension $\ell\ll n^{d}$, then the cost of forming $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is only $O(\ell dnk+\ell k^{d})$, suppressing log factors, instead of $O(\ell n^{d}k^{d})$ for naive multiplication. This is accomplished by sketching each factor matrix with independent CountSketch operators, then convolving with Fast Fourier Transforms (FFTs).
@@
To that end, let $h_{j}:[n]\rightarrow[\ell]$ for $j=1,\ldots,d$ be 3-wise independent hash functions, i.e. for any distinct $i_{1},i_{2},i_{3}\in[n]$, the hash codes $h_{j}(i_{1})$, $h_{j}(i_{2})$, $h_{j}(i_{3})$ are independent random variables, and $h_{j}(i)$ is uniformly distributed in $\ell$] for any fixed $i\in[n]$. Let $s_{j}:[n]\rightarrow\{+1,-1\}$ be 4-wise independent sign functions (so that analogously, $s_{j}(i_{1})$, $s_{j}(i_{2})$, $s_{j}(i_{3})$, $s_{j}(i_{4})$ are independent random variables with $s_{j}(i)\sim\textup{Uniform}\{\pm 1\}$. We form independent CountSketch matrices $\bm{\mathsf{S}}^{(j)}=\bm{\mathsf{P}}^{(j)}\bm{\mathsf{D}}^{(j)}$ for $j=1,\ldots,d$, using the hash functions $h_{j}$ to form $\bm{\mathsf{P}}^{(j)}$ and $s_{j}(i)=\bm{\mathsf{D}}^{(j)}_{i,i}$ for $i\in[n]$. The application of $\bm{\mathsf{S}}^{(j)}$ to each column $\bm{\mathsf{A}}^{(j)}_{:,r_{j}}$, for $r_{j}=1,\ldots,k$, can be represented by an $(\ell-1)$-degree polynomial
@@
and we define $\bm{\mathsf{c}}_{r_{j}}^{(j)}$ to be the vector of coefficients $(c^{(j)}_{1,r_{j}},\ldots,c^{(j)}_{\ell,r_{j}})$, which are grouped in (57) by hash value.
Let $\bm{\mathsf{T}}$ be the TensorSketch operator, which is defined as the CountSketch operator formed from hash function $H$ and sign function $S$ given by
@@
Then the $r$th column of $\bm{\mathsf{T}}(\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)})=\bm{\mathsf{T}}\bm{\mathsf{Z}}$ for $r=1,\ldots,k^{d}$ can also be represented as a polynomial
@@
where * denotes the Hadamard product, and $i$ in (60) corresponds to the tuple $(i_{1},\ldots,i_{d})\in[n]^{d}$, and similarly $r$ corresponds to the tuple $(r_{1},\ldots,r_{d})\in[k]^{d}$. In other words,
$\bm{\mathsf{T}}\bm{\mathsf{Z}}=\textup{FFT}^{-1}\left(\left({\bigodot}_{j=1}^{d}\left(\textup{FFT}\left(\bm{\mathsf{S}}^{(j)}\bm{\mathsf{A}}^{(j)}\right)\right)^{\top}\right)^{\top}\right),$ (63)
@@
Using TensorSketch in HOOI We now turn our attention back to the randomized HOOI algorithm of Malik2018, which uses TensorSketch to reduce the size of each least squares subproblem in the two primary steps of HOOI listed above. The least squares problem to find the factor matrix for each mode $j=1,\ldots,d$ is highly overdetermined, involving a Kronecker product $\bm{\mathsf{Z}}^{(j)}$ of size $N^{(-j)}\times(\prod_{k\neq j}R_{k})$. The least squares problem to find the core tensor is also highly overdetermined, involving a Kronecker product of size $\prod_{j=1}^{d}N_{j}\times\prod_{j=1}^{d}R_{j}$, or $n^{d}\times k^{d}$ for convenience. To compute the factor matrices and core tensor, we form independent TensorSketches to solve smaller sketched least squares problems corresponding to both of these steps. Pseudocode for R-HOOI, equivalently (Malik2018, TUCKER-TS), is given in Algorithm 10. This algorithm can be implemented as a single-pass algorithm through the input tensor, cf. (Malik2018, Algorithm S3).
The dominant cost of Algorithm 10 (single-pass) is $O(dnk^{2}+\ell dk^{d}+\ell_{c}k^{d}+k^{2d})$, or roughly $k^{O(d)}$, versus the dominant cost of $O(n^{d})$ for deterministic HOOI (Algorithm 7), if we suppress $\log$ factors and assume sketching dimension $\ell=\ell_{1}=\cdots=\ell_{d}$. We refer the reader to (Malik2018, S3.2.6) for details on the complexity analysis. However, we observe that the single-pass R-HOOI algorithm as presented in Malik2018 uses the same TensorSketches for sketching the updated Kronecker products until some convergence criteria is met. In practice, the performance of R-HOOI as written in Algorithm 10 is curiously better in terms of error reduction than if new independent TensorSketch operators were drawn in every iteration, though it is unclear why this is the case. However, for the theoretical guarantees of DiaoSongSunWoodruff2018 to hold for TensorSketched least squares, it is technically necessary to form new independent TensorSketch operators in each iteration of R-HOOI. In Malik2018, the numerical results suggest that sketching dimensions $\ell=Ck^{d-1}$ and $\ell_{c}=Cr^{d}$ for constant $C>4$ (e.g., $C=10$) work well in practice.
Algorithm 10 Randomized Higher-order Orthogonal Iteration with TensorSketch (R-HOOI)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, factor matrix target ranks $R_{1},\ldots,R_{d}$, sketching dimensions $\ell_{1},\ldots,\ell_{d}$, $\ell_{c}$
4:Draw TensorSketch matrices $\bm{\mathsf{T}}^{(j)}\in\mathbb{R}^{\ell_{j}\times(\prod_{i\neq j}N_{i})}$ for $i,j\in[d]$ and $\bm{\mathsf{T}}^{(c)}\in\mathbb{R}^{\ell_{c}\times\prod_{j}N_{j}}$
5:while not converged do
6:  for $j=1,\ldots,d$ do
7:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{T}}^{(j)}\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{T}}^{(j)}\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
8:  end for
9:  $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bm{\mathsf{T}}^{(c)}\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{T}}^{(c)}\bm{\mathsf{x}}\right\|^{2}_{2}$ $\triangleright$ Column vectorization $\bm{\mathsf{x}}$ of $\bm{\mathcal{X}}$
10:  Re-fold vectorization $\bm{\mathsf{g}}$ into $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$
11:end while
@@
We end this section by briefly highlighting the recent work of Hashemi2025 on Randomized Tucker decompositions with Single-Mode Sketching (RTSMS). The RTSMS algorithm sketches the input tensor one mode at a time, i.e. sketching on the left of the mode-$j$ unfolding vs. sketching on the right, to avoid the computational bottleneck of multiplying along the larger dimension $N^{(-j)}=\prod_{i\neq j}N_{i}$. The algorithm is also single-pass, so that the input tensor is not revisited after the first sketch, and rank-adaptive if the target ranks for the factor matrices are unknown a priori.
To illustrate, suppose we want to compute a rank-$(R_{1},R_{2},R_{3})$ HOSVD that approximates an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$. (We omit rank-adaptivity and iterative refinement; the reader is referred to Hashemi2025 for practical implementation details.) We initialize tensors
$\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{X}}\ \textup{and}\ \bm{\mathcal{G}}^{\textup{new}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}^{(1)}\iff\bm{\mathsf{G}}^{\textup{new}}_{(1)}=\left(\bm{\mathsf{\Omega}}^{(1)}\right)^{\top}\bm{\mathsf{X}}_{(1)}\in\mathbb{F}^{\tilde{R_{1}}\times(N_{2}N_{3})}$
for some sketching dimension $\tilde{R_{1}}>R_{1}$. There exists a factor matrix $\bm{\mathsf{A}}^{(1)}\in\mathbb{F}^{N_{1}\times\tilde{R_{1}}}$ such that $\bm{\mathcal{G}}^{\textup{new}}\approx\bm{\mathcal{G}}^{\textup{old}}\times_{1}\bm{\mathsf{A}}^{(1)}$. In practice, this is computed as an approximate solution of the corresponding least squares problem, which is being sub-sampled from the right, cf. (Hashemi2025, Section 4.2). We then update $\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{G}}^{\textup{new}}$.
For the next mode, we first update $\bm{\mathcal{G}}^{\textup{new}}$ via mode-2 sketching:
$\bm{\mathsf{G}}^{\textup{new}}_{(2)}=\left(\bm{\mathsf{\Omega}}^{(2)}\right)^{\top}\bm{\mathsf{G}}_{(2)}^{\textup{old}}\in\mathbb{F}^{\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})}$
@@
The computational complexity of RTSMS is dominated by the first TTM $\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}_{1}$. Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the dominant cost is $O(kn^{d})$ to compute the first TTM, but the sketch size is only $k\times n$ vs. $k\times n^{d-1}$ for R-HOSVD. The least squares solves are also significantly reduced by sub-sampling and by computing a generalized Nyström decomposition vs. SVD for the factor matrices. The full tensor does not need to be sketched again in modes $j=2,\ldots,d$, and the core tensor is available immediately upon termination, as in the ST-HOSVD algorithm. Additional details can be found in Hashemi2025.
@@ 6.3 Randomized Algorithms for Tensor ID/CUR Decompositions
As we observed in Sect. 5.3, the tensor ID/CUR decomposition has been formulated in many ways in the literature, with no one standard representation. However, there exists a natural categorization of tensor ID/CUR decompositions, based on whether the factor matrices or core tensors in the computed decompositions are comprised of entries of the input tensor. In this section, we explore how randomization can be leveraged in prototypical algorithms to compute a Tucker ID/CUR decomposition from each category. We also discuss recent work on randomized algorithms for tensor ID/CUR in the CP format, which can be viewed as a form of input tensor pruning.
@@
We first consider randomized algorithms for tensor ID/CUR factorizations in the Tucker format of the first category, where the core tensor in the compressed representation is a sub-tensor of the input tensor. In particular, we focus on the tensor ID/CUR decompositions that were recently characterized in Cai2024; Cai2021, termed the “Chidori” and “Fiber” tensor CUR factorizations. This work was inspired by numerous investigations of randomized sampling algorithms for tensor CUR decompositions, including CaiafaCichocki2009; Caiafa2010; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
In Cai2021, given an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the rank-$(R_{1},\ldots,R_{d})$ tensor CUR factorization is defined as
@@
for index sets $I_{j}\subset[N_{j}]$, $j=1,\ldots,d$. The mode-$j$ factor matrix $\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger}$ is defined in terms of the mode-$j$ unfolding:
@@
where $J_{j}\subset[N^{(-j)}]$. If indices $J_{j}$ are computed independently from indices $I_{j}$, the decomposition in (64) is called a “Fiber” CUR decomposition in Cai2021. If $J_{j}=\otimes_{k\neq j}I_{k}$, then (64) is called a “Chidori” CUR decomposition. By (Cai2021, Theorem 3.3), (HammHuang2020, Corollary 5.2), if the indices $I_{j}$ and $J_{j}$ are each sampled uniformly with $|I_{j}|=O(R_{j}\log N_{j})$ and $|J_{j}|=O(R_{j}\log N^{(-j)})$, then the computed Tucker decomposition is a rank-$(R_{1},\ldots,R_{d})$ approximation with high probability. We present pseudocode for either the Chidori or Fiber tensor CUR decomposition in Algorithm 11.
With uniform sampling, the computational complexity of computing either tensor CUR decomposition is dominated by the cost of the pseudoinverse of $\bm{\mathsf{U}}_{j}$, which is $O(k^{d+1}\log^{d}(n))$ for Chidori and $O(k^{2}\log^{2}(n))$ for Fiber, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$ as before. Recall that the HOSVD has complexity $O(kn^{d})$ by comparison.
Algorithm 11 Randomized Tucker ID/CUR with interpolating core
[…]
2:chidori or fiber decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}(\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger})$
3:for $j=1,\ldots,d$ do
4:  if chidori then
5:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
6:   Set $J_{j}=\otimes_{i\neq j}I_{i}$
7:  end if
8:  if fiber then
9:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
10:   Sample indices $J_{j}$ from $N^{(-j)}]$, $|J_{j}|=T_{j}$ without replacement according to $\bm{\mathsf{q}}_{j}$
11:  end if
12:  Form $\bm{\mathsf{C}}_{j}=\bm{\mathsf{X}}_{(j)}(:,J_{j})$
13:  Form $\bm{\mathsf{U}}_{j}=\bm{\mathsf{C}}_{j}(I_{j},:)$
14:end for
[…]
The recent work of Zhang2025 proposes a similar factorization to (64), called the “CoreID.” The CoreID algorithm is reminiscent of the sequentially-truncated HOSVD, specifically the “structure-preserving” ST-HOSVD developed for sparse input tensors in MinsterSaibabaKilmer2020. Namely, in the CoreID, as each mode’s factor matrix is computed, the input tensor is compressed along that mode before proceeding to the next one, so that each subsequent unfolding is reduced in size.
Through matricization, computing the CoreID is reduced to computing a randomized matrix row ID in each mode, cf. Algorithm 4. For $j=1,\ldots,d$, the mode-$j$ unfolding of the tensor is sketched by a random matrix $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}_{j}.$ Then a row ID of $\bm{\mathsf{Y}}_{j}$ is computed by some skeleton selection method, e.g., greedy pivoting or squared-norm sampling. The selected skeletons $I_{j}$ form the core sub-tensor, and the mode-$j$ factor matrix for the CoreID is defined as the associated interpolation matrix. The tensor is then compressed along mode $j$ as in ST-HOSVD before proceeding to the next mode. Sketching is the dominating cost in this algorithm, which is why many of the numerical investigations in Zhang2025 consider structured input tensors for fast sketching, e.g., a sparse input tensor or a CP-factorized input tensor.
@@
We now turn our attention to randomized algorithms for tensor ID/CUR factorizations in the Tucker format in which the factor matrices contain entries of the input tensor. We illustrate such factorizations with the “higher-order ID” of Saibaba2016, which we contrast with the “SatID” of Zhang2025.

### v2

@@ 6.2 Randomized Algorithms for Tucker Decompositions
The CountSketch operator $\bm{\mathsf{S}}$ can be rapidly applied to a matrix $\bm{\mathsf{A}}$ in $O(\textup{nnz}(\bm{\mathsf{A}}))$ operations, without explicitly forming $\bm{\mathsf{S}}$ or $\bm{\mathsf{A}}$. If $\bm{\mathsf{S}}\in\mathbb{R}^{J\times I}$ and $\bm{\mathsf{A}}\in\mathbb{F}^{I\times L}$, typically $J\ll I$, we form $\bm{\mathsf{Y}}=\bm{\mathsf{S}}\bm{\mathsf{A}}$ by hashing each row $i$ of $\bm{\mathsf{A}}$ with an integer $h$ sampled uniformly from $[J]$, assigning it a value $s$ of $\pm 1$ with equal probability, then summing the rows with the same hash value (i.e. adding $s\bm{\mathsf{A}}_{i,:}$ to the $h$th row of $\bm{\mathsf{Y}}$ if $\bm{\mathsf{A}}_{i,:}$ has hash value $h$).
@@
A TensorSketch operator is a specific type of CountSketch operator, which can be rapidly applied to Kronecker products, e.g., $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)}\in\mathbb{F}^{n^{d}\times k^{d}}$. If $\bm{\mathsf{T}}:\mathbb{R}^{n^{d}}\rightarrow\mathbb{R}^{\ell}$ is a TensorSketch operator, then the cost of sketching $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is $O(\ell dnk+\ell k^{d})$, suppressing log factors, instead of $O(\ell n^{d}k^{d})$ for naive multiplication. This is accomplished by sketching each factor matrix with independent CountSketch operators, then convolving with Fast Fourier Transforms (FFTs).
@@
Let $h_{j}:[n]\rightarrow[\ell]$ for $j=1,\ldots,d$ be 3-wise independent hash functions, i.e. for any distinct $i_{1},i_{2},i_{3}\in[n]$, the hash codes $h_{j}(i_{1})$, $h_{j}(i_{2})$, $h_{j}(i_{3})$ are $\textup{Uniform}\{[\ell]\}$. Let $s_{j}:[n]\rightarrow\{+1,-1\}$ be 4-wise independent sign functions (i.e. $\textup{Uniform}\{\pm 1\}$). We form independent CountSketch matrices $\bm{\mathsf{S}}^{(j)}=\bm{\mathsf{P}}^{(j)}\bm{\mathsf{D}}^{(j)}$ for $j=1,\ldots,d$, using the hash functions $h_{j}$ to form $\bm{\mathsf{P}}^{(j)}$ and $s_{j}(i)=\bm{\mathsf{D}}^{(j)}_{i,i}$ for $i\in[n]$. Applying $\bm{\mathsf{S}}^{(j)}$ to each column $\bm{\mathsf{A}}^{(j)}_{:,r_{j}}$, for $r_{j}=1,\ldots,k$, can be represented by an $(\ell-1)$-degree polynomial
@@
where $\bm{\mathsf{c}}_{r_{j}}^{(j)}$ are coefficients $(c^{(j)}_{1,r_{j}},\ldots,c^{(j)}_{\ell,r_{j}})$, grouped in (50) by hash value.
The TensorSketch operator $\bm{\mathsf{T}}$ is defined as the CountSketch operator comprised of hash function $H$ and sign function $S$ given by
@@
The $r$-th column of $\bm{\mathsf{T}}(\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)})=\bm{\mathsf{T}}\bm{\mathsf{Z}}$, $r=1,\ldots,k^{d}$, is the polynomial
@@
where * denotes the Hadamard product, $i$ in (53) corresponds to the tuple $(i_{1},\ldots,i_{d})\in[n]^{d}$, and $r$ corresponds to the tuple $(r_{1},\ldots,r_{d})\in[k]^{d}$. In other words,
$\bm{\mathsf{T}}\bm{\mathsf{Z}}=\textup{FFT}^{-1}\left(\left({\bigodot}_{j=1}^{d}\left(\textup{FFT}\left(\bm{\mathsf{S}}^{(j)}\bm{\mathsf{A}}^{(j)}\right)\right)^{*}\right)^{*}\right),$ (56)
@@
Using TensorSketch in HOOI We now turn our attention back to the randomized HOOI algorithm of Malik2018, called TUCKER-TS, which uses TensorSketch in each least squares subproblem in the two main steps above. In Step 1, the factor matrix sub-problems for modes $j=1,\ldots,d$ are highly overdetermined, involving an $n^{d-1}\times k^{d-1}$ Kronecker product $\bm{\mathsf{Z}}^{(j)}$. The core tensor sub-problem in Step 2 is also highly overdetermined, involving a $n^{d}\times k^{d}$ Kronecker product. For both steps, we form independent TensorSketches to solve smaller sketched least squares problems, as in Sect. 3.3. TUCKER-TS can also be made single-pass; cf. (Malik2018, Algorithm S3).
The dominant cost of randomized HOOI with TensorSketch (TUCKER-TS) is $O(dnk^{2}+\ell dk^{d}+\ell_{c}k^{d}+k^{2d})$, or roughly $k^{O(d)}$, versus the dominant cost of $O(n^{d})$ for classical HOOI (Algorithm 7), suppressing $\log$ factors and letting $\ell=\ell_{1}=\cdots=\ell_{d}$. See (Malik2018, S3.2.6) for details on the complexity analysis. We note that the single-pass algorithm presented in Malik2018 uses the same TensorSketch operators on the updated Kronecker products until some convergence criteria is met. For the theoretical guarantees of DiaoSongSunWoodruff2018 to hold, it is technically necessary to form new independent TensorSketch operators in each iteration; however, in practice, it is noted in Malik2018 that re-using TensorSketch operators is more efficient and works just as well. Their numerical results suggest $\ell=Ck^{d-1}$ and $\ell_{c}=Cr^{d}$ for constant $C>4$ (e.g., $C=10$) are usually appropriate.
@@
We end this section by highlighting the recent work of Hashemi2025 on Randomized Tucker decompositions with Single-Mode Sketching (RTSMS). The RTSMS algorithm sketches the input tensor one mode at a time, i.e. sketching on the left of the mode-$j$ unfolding vs. sketching on the right, to avoid the computational bottleneck of multiplying along the larger dimension $N^{(-j)}\equiv n^{d-1}$. The algorithm is also single-pass and rank-adaptive.
To illustrate, suppose we want to compute a rank-$(R_{1},R_{2},R_{3})$ HOSVD approximating $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$. (We omit rank-adaptivity and iterative refinement; see Hashemi2025 for practical implementation details.) We initialize tensors
$\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{X}}\ \textup{and}\ \bm{\mathcal{G}}^{\textup{new}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}^{(1)}\iff\bm{\mathsf{G}}^{\textup{new}}_{(1)}=\left(\bm{\mathsf{\Omega}}^{(1)}\right)^{*}\bm{\mathsf{X}}_{(1)}\in\mathbb{F}^{\tilde{R_{1}}\times(N_{2}N_{3})}$
for some sketching dimension $\tilde{R_{1}}>R_{1}$. There exists a factor matrix $\bm{\mathsf{A}}^{(1)}\in\mathbb{F}^{N_{1}\times\tilde{R_{1}}}$ such that $\bm{\mathcal{G}}^{\textup{new}}\approx\bm{\mathcal{G}}^{\textup{old}}\times_{1}\bm{\mathsf{A}}^{(1)}$. In practice, this is computed as an approximate solution of the corresponding least squares problem, which is being sub-sampled from the right, cf. (Hashemi2025, Section 4.2). We update $\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{G}}^{\textup{new}}$, and for the next mode, update $\bm{\mathcal{G}}^{\textup{new}}$ via
$\bm{\mathsf{G}}^{\textup{new}}_{(2)}=\left(\bm{\mathsf{\Omega}}^{(2)}\right)^{*}\bm{\mathsf{G}}_{(2)}^{\textup{old}}\in\mathbb{F}^{\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})}$
@@
The computational complexity of RTSMS is dominated by the first TTM $\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}_{1}$. The dominant cost is $O(kn^{d})$ to compute the first TTM, but the sketch size is only $k\times n$ vs. $k\times n^{d-1}$ for R-HOSVD. The least squares solves are also significantly reduced by sub-sampling and by computing a generalized Nyström decomposition vs. SVD. The full tensor does not need to be sketched again in later modes, and the core tensor is available immediately upon termination. Additional details can be found in Hashemi2025.
@@ 6.3 Randomized Algorithms for Tensor ID/CUR Decompositions
As we observed in Sect. 5.3, the tensor ID/CUR decomposition has been formulated in many ways in the literature, with no one standard representation. However, there exists a natural categorization of tensor ID/CUR decompositions, based on whether the factor matrices or core tensors in the computed decompositions are comprised of entries of the input tensor. We discuss randomized algorithms to compute a Tucker ID/CUR decomposition from each category, as well as recent work on CP tensor ID/CUR.
@@
We first consider randomized algorithms for Tucker tensor ID/CUR factorizations of the first type, where the core tensor is a sub-tensor of the input tensor. In particular, we focus on the tensor ID/CUR decompositions that were recently characterized in Cai2021, termed the “Chidori” and “Fiber” tensor CUR factorizations. This work was inspired by numerous investigations of tensor CUR decompositions, including Caiafa2010; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
In Cai2021, given $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the rank-$(R_{1},\ldots,R_{d})$ tensor CUR factorization is
@@
for index sets $I_{j}\subset[N_{j}]$, $j=1,\ldots,d$. The mode-$j$ factor matrix $\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger}$ is defined by
@@
where $J_{j}\subset[N^{(-j)}]$. If indices $J_{j}$ are computed independently from indices $I_{j}$, the decomposition in (57) is called a “Fiber” CUR decomposition in Cai2021. If $J_{j}=\otimes_{k\neq j}I_{k}$, then (57) is called a “Chidori” CUR decomposition.
By (Cai2021, Theorem 3.3), (HammHuang2020, Corollary 5.2), if the indices $I_{j}$ and $J_{j}$ are each sampled uniformly with $|I_{j}|=O(R_{j}\log N_{j})$ and $|J_{j}|=O(R_{j}\log N^{(-j)})$, then the computed Tucker decomposition is a rank-$(R_{1},\ldots,R_{d})$ approximation with high probability, under certain incoherence assumptions on the input tensor; cf. Dong2021TensorCompletion. With uniform sampling, the computational complexity of computing either CUR decomposition is dominated by the cost of $\bm{\mathsf{U}}_{j}^{\dagger}$, which is $O(k^{d+1}\log^{d}(n))$ for Chidori and $O(k^{2}\log^{2}(n))$ for Fiber, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$ as before.
The recent work of Zhang2025 proposes a similar factorization to (57), called the “CoreID.” The CoreID algorithm is reminiscent of the “structure-preserving” ST-HOSVD developed for sparse input tensors in MinsterSaibabaKilmer2020. Namely, in the CoreID, as each factor matrix is computed, the input tensor is compressed along that mode before proceeding to the next one, so that each subsequent unfolding is reduced in size. Through matricization, computing the CoreID is reduced to computing a randomized matrix row ID in each mode, cf. Algorithm 4. For $j=1,\ldots,d$, the mode-$j$ unfolding of the tensor is sketched by a random matrix $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}_{j}.$ Then a row ID of $\bm{\mathsf{Y}}_{j}$ is computed by some skeleton selection method. The selected skeletons $I_{j}$ form the core sub-tensor, and the mode-$j$ factor matrix for the CoreID is defined as the associated interpolation matrix. The tensor is then compressed along mode $j$ before proceeding to the next. Sketching is the dominant cost, so many of the numerical experiments in Zhang2025 consider structured input tensors for fast sketching, e.g., a sparse input tensor or a CP-factorized input tensor.
@@
We now turn our attention to randomized algorithms for tensor ID/CUR factorizations in the Tucker format in which the factor matrices contain entries of the input tensor. We detail the “higher-order ID” of Saibaba2016, contrasted with the “SatID” of Zhang2025.